A modified​ box-and-whisker plot is a​ box-and-whisker plot that uses symbols to identify outliers. The horizontal line of a modified​ box-and-whisker plot extends as far as the minimum data entry that is not an outlier and the maximum data entry that is not an outlier.​ (a) Identify any outliers and​ (b) draw a modified​ box-and-whisker plot that represents the data set.

11,2,8,13,8,13,10,10,12,14,11,11,24,16,10

​(a) Identify any outliers. Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.
Answer #1

Median=0.5(n+1)th value =    8th   value of sorted data
=   11.000  
      
quartile , Q1 = 0.25(n+1)th value=   4th   value of sorted data
=   10  
      
Quartile , Q3 = 0.75(n+1)th value=   12th   value of sorted data
=   13  
      
IQR = Q3-Q1 =    3  
      
1.5IQR =    4.5  
      
lower bound=Q1-1.5IQR=   5.5  
      
upper bound=Q3+1.5IQR=   17.5  
      
outlier =values outside lower bound and upper bound      
total outlier below lower bound=   1  
total outlier above upper bound=   1  
total outlier =    2  
outliers are 2 and 24